### v2

@@ 1 Background
After a close examination of the SLQ method, it turns out that Ubaru, Chen and Saad’s result works for the cases when the quadrature nodes are symmetric [33]. Whereas Cortinovis and Kressner pointed out that quadrature nodes generated in the SLQ process may be asymmetric [6]. We propose Theorem 3.6 that guarantees the symmetric quadrature nodes under certain assumptions, but perhaps the asymmetric cases arise more often in practice. For instance, Fig. 1 illustrates that the quadrature nodes generated by SLQ are asymmetric based on the matrix nd3k from [7]. Another point worth noticing is that the SLQ method is based on a quadrature rule corresponding to a discontinuous measure defined by the spectrum of the underlying matrix (see (5) below). Unlike the continuous measure case, under such a discontinuous measure the quadrature error will remain unchanged when an affine transformation from the reference domain $\left[-1,1\right]$ to the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$ is applied (this will be detailed in Section 3.4). With consideration of these two points, this paper updates estimation on the query numbers and the Lanczos steps based on the asymmetric quadrature rule, and further develops a probabilistic relative error bound for log determinant estimation. The main results are presented in Theorem 3.16, Theorem 3.18 and Theorem 4.1.
@@
Figure 1: Illustration of asymmetric quadrature nodes generated for the nd3k matrix from [7]. The initial normalized Rademacher vector $\boldsymbol{v}$ is chosen as $(1,\ldots,1,-1,\ldots,-1)^{T}/\sqrt{9000}\in\mathbb{R}^{9000}$, where there are equal numbers of 1s and -1s. Fig. 1a and Fig. 1c show quadrature nodes (eigenvalues) in the physical domain $\left[\theta_{1},\theta_{10}\right]$ and $\left[\theta_{1},\theta_{100}\right]$, while Fig. 1b and Fig. 1d depict the quadrature nodes that are transferred to the reference interval $[-1,1]$.
@@ 2 SLQ basics
$\log\det A=\log\prod_{j=1}^{n}\lambda_{j}=\sum_{j=1}^{n}\log\lambda_{j}=\operatorname{tr}(\log(A)).$ (1)
@@
$\operatorname{tr}(\log(A))\approx\operatorname{tr}_{N}(\log(A))=\frac{1}{N}\sum_{i=1}^{N}{\boldsymbol{z}^{(i)}}^{T}\log(A)\boldsymbol{z}^{(i)}=\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{v}^{(i)}}^{T}\log(A)\boldsymbol{v}^{(i)},$ (2)
@@
$=\frac{n}{N}\sum_{i=1}^{N}\ {\boldsymbol{\mu}^{(i)}}^{T}\log(\Lambda)\boldsymbol{\mu}^{(i)}.$ (3)
[…]
$\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{\mu}^{(i)}}^{T}\log(\Lambda)\boldsymbol{\mu}^{(i)}$ $=\frac{n}{N}\sum_{i=1}^{N}\sum_{j=1}^{n}\log(\lambda_{j}^{(i)})\left[{\mu_{j}^{(i)}}\right]^{2}$ (4)
@@
$\mu^{(i)}(t)=\left\{\begin{aligned} &\quad 0,&{\rm if}\ t<\lambda_{1}=a,\\ &\sum_{j=1}^{k-1}\left[{\mu_{j}^{(i)}}\right]^{2},&{\rm if}\ \lambda_{k-1}\leq t<\lambda_{k},k=2,...,n,\\ &\sum_{j=1}^{n}\left[{\mu_{j}^{(i)}}\right]^{2},&{\rm if}\ t\geq\lambda_{n}=b.\end{aligned}\right.$ (5)
@@
$\frac{n}{N}\sum_{i=1}^{N}I^{(i)}\approx\frac{n}{N}\sum_{i=1}^{N}I_{m}^{(i)}=\frac{n}{N}\sum_{i=1}^{N}\sum_{k=1}^{m+1}\tau_{k}^{(i)}\log(\theta_{k}^{(i)}),$ (6)
where $m$ is the Lanczos step, and $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$, $\{\theta_{k}^{(i)}\}_{k=1}^{m+1}$ are the quadrature weights and nodes. According to [13], the weights $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$ are computed by the squares of the $1^{st}$ elements of the normalized eigenvectors $\{\boldsymbol{y}_{k}^{(i)}\}_{k=1}^{m+1}$ of $T_{m+1}^{(i)}$, and the nodes (also known as the Ritz values) $\{\theta_{k}^{(i)}\}_{k=1}^{m+1}$ are the corresponding eigenvalues $\{\lambda_{k}\}_{k=1}^{m+1}$ of $T_{m+1}^{(i)}$. Algorithm 1 [12, Section 4.1] outlines how to compute $m+1$ Lanczos quadrature nodes and weights given a real symmetric matrix and a randomly generated Rademacher vector.
@@
10:Calculate the eigenpairs of $T_{m+1}$: $[V,D]=\texttt{eig}(T_{m+1})$.
@@
Algorithm 2 [33, Section 3] summarizes how the SLQ method approximates the log determinant for any SPD matrix.
@@
3:     Apply $m$-step Algorithm 1 to $A$ with the initial vector $\boldsymbol{v}^{(i)}$ and obtain $\{\theta^{(i)}_{k}\}_{k=1}^{m+1}$ and $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$.
@@ 3 Error Analysis
According to [33, p. 9-10], the computational cost of stochastic Lanczos quadrature is $\mathcal{O}((\#nz(A)+nm^{2})N)$, where #$\#nz(A)$ denotes the number of non-zeros in symmetric matrix $A$ and $\mathcal{O}(nm^{2})$ represents the orthogonalization cost spent in the Lanczos iteration. Normally, the steps of Lanczos iteration $m$ and the number of Rademacher distributed vectors $N$ are expected to be much smaller than the matrix size $n$.
@@ 3.1 Error for symmetric quadrature nodes
[33, Theorem 4.2] Let $f$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci ±$\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $(m+1)$-point Lanczos quadrature approximation satisfies
$|I-I_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-2}}\rho^{-2m-2}.$ (7)
@@
See [33, Section 4.1]. It should be pointed out that the result of Theorem 3.1 assumes that the quadrature nodes are symmetric.
@@
[33, Corollary 4.5] Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
@@ 3.2 Existence of symmetric quadrature nodes
As previously mentioned, Theorem 3.1 and Corollary 3.3 are valid when the Gauss quadrature nodes are symmetric, but when do these nodes preserve this property? In order to answer this question, we propose the following Theorem 3.6 with the aid of Definition 3.5.
@@
$A=\begin{bmatrix}A_{1,1}&\cdots&A_{1,q}\\ \vdots&&\vdots\\ A_{q,1}&\cdots&A_{q,q}\end{bmatrix},$ (8)
@@
$\begin{bmatrix}D_{1}&F\\ E&D_{2}\end{bmatrix},$ (9)
@@
$C=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix},$ (10)
where $O$ denotes matrix of zeros and $B\in\mathbb{C}^{m\times n}$. Such matrices have symmetric eigenvalue distribution about 00, proved by Theorem A.1 in Appendix A [21, 3]. Here we give a more concise theorem that proves the symmetry of eigenvalues when $m=n$.
@@
has a symmetric eigenvalue distribution. Precisely, $\left(\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}\right)$ and $\left(-\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}\right),\\ i=1,\cdots,n$ are the eigenpairs of $C$.
@@
Now we have checked the existence of matrices with symmetric eigenvalue distribution. Furthermore, when the conditions in Theorem 3.6 hold, the Jacobi matrix $T_{m+1}$ generated by $m$-step Lanczos iteration has constant diagonal entries.
@@
$\begin{bmatrix}\bar{\lambda}I_{m}&B\\ B^{H}&\bar{\lambda}I_{n}\end{bmatrix}.$ (11)
Thus, based on Corollary A.3, the Jacobi matrix $T_{m+1}$ generated by $m$-step Lanczos iteration have symmetric eigenvalue distribution, indicating that the $(m+1)$ quadrature nodes are symmetric about $\bar{\lambda}$. In addition, one may find several tests on Theorem 3.6 in Appendix B.
@@ 3.3 Error for asymmetric quadrature nodes
Though Theorem 3.6 gives sufficient conditions for symmetric quadrature nodes, the conditions are not always satisfied in practice. As shown in Section 1, Fig. 1 demonstrates that Lanczos quadrature nodes might be asymmetric. Therefore, Theorem 3.12 derives an error estimation for the Lanczos quadrature with asymmetric quadrature nodes.
@@
[6, Corollary 3] Let $g$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci ±$\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $(m+1)$-point Lanczos quadrature approximation with asymmetric quadrature nodes satisfies
$|I-I_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-1}}\rho^{-2m-2}.$ (12)
@@
The proofs of Theorem 3.12 and Theorem 3.1 differ by a single step, namely the calculation of an infinite geometric progression
$|I-I_{m}|\leq\sum_{j=2m+2}^{\infty}\frac{4M_{\rho}}{\rho^{j}}.$ (13)
@@
$\sum_{j=2m+2}^{\infty}\frac{4M_{\rho}}{\rho^{j}}=\frac{4M_{\rho}}{\rho^{2m+2}}\cdot\frac{1}{1-\rho^{-1}}=\frac{4M_{\rho}}{1-\rho^{-1}}\rho^{-2m-2}.$ (14)
@@
Theorem 3.12 is in fact [6, Corollary 3] but with a redefinition of the parameter $m$. In [6, Algorithm 1], $m$ corresponds to $m$ quadrature nodes. Whereas, in Theorem 3.12, $m$ corresponds to $(m+1)$ quadrature nodes, so does [33].
@@ 3.4 Error under affine transform
Theorem 3.12 assumes that function $g$ is analytic in the reference domain $\left[-1,1\right]$ and then gives an error estimation in the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$. In order to conduct analysis back on the physical interval, a proper affine transform should be included. The SLQ method is based on a quadrature rule corresponding to a discontinuous measure (5). Unlike the continuous measure case, under such a discontinuous measure the quadrature error will remain unchanged when using an affine transformation from the reference domain $\left[-1,1\right]$ to the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$ [16]. Precisely, consider the integrals
[…]
@@
[…]
where $\mu^{(i)}$ are defined in (5), $h$ is an affine transform that shifts the interval $[-1,1]$ to $[\lambda_{\min},\lambda_{\max}]$, $g=f\circ h$, and $g$ satisfies the assumptions in Theorem 3.12. These integrals are in fact the finite sums of the form
[…]
@@
The measure in (15) does not satisfy assumption (16) since $d\mu^{(i)}(h(t))$ assigns positive measure $\left[\mu_{j}^{(i)}\right]^{2}$ on single point $h^{-1}(\lambda_{j})$, while the Lebesgue measure $dt$ assigns zero on any single point. This may be the reason that (15) fails.

### v3

@@ 1 Background
After a close examination of the SLQ method, it turns out that Ubaru, Chen and Saad’s result works for the cases when the quadrature nodes are symmetric [33]. Whereas Cortinovis and Kressner pointed out that quadrature nodes generated in the SLQ process may be asymmetric [6]. We propose Theorem˜3.6 that guarantees the symmetric quadrature nodes under certain assumptions, but perhaps the asymmetric cases arise more often in practice. For instance, Fig.˜1 illustrates that the quadrature nodes generated by SLQ are asymmetric based on the matrix nd3k from [7]. Another point worth noticing is that the SLQ method is based on a quadrature rule corresponding to a discontinuous measure defined by the spectrum of the underlying matrix (see (5) below). Unlike the continuous measure case, under such a discontinuous measure the quadrature error will remain unchanged when an affine transformation from the reference domain $\left[-1,1\right]$ to the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$ is applied (this will be detailed in Section˜3.4). With consideration of these two points, this paper updates estimation on the query numbers and the Lanczos steps based on the asymmetric quadrature rule, and further develops a probabilistic relative error bound for log determinant estimation. The main results are presented in Theorem˜3.16, Theorem˜3.18 and Theorem˜4.1.
@@
Figure 1: Illustration of asymmetric quadrature nodes generated for the nd3k matrix from [7]. The initial normalized Rademacher vector $\boldsymbol{v}$ is chosen as $(1,\ldots,1,-1,\ldots,-1)^{T}/\sqrt{9000}\in\mathbb{R}^{9000}$, where there are equal numbers of 1s and -1s. Fig.˜1a and Fig.˜1c show quadrature nodes (eigenvalues) in the physical domain $\left[\theta_{1},\theta_{10}\right]$ and $\left[\theta_{1},\theta_{100}\right]$, while Fig.˜1b and Fig.˜1d depict the quadrature nodes that are transferred to the reference interval [1,1][-1,1].
@@ 2 SLQ basics
(1) $\log\det A=\log\prod_{j=1}^{n}\lambda_{j}=\sum_{j=1}^{n}\log\lambda_{j}=\operatorname{tr}(\log(A)).$
@@
(2) $\operatorname{tr}(\log(A))\approx\operatorname{tr}_{N}(\log(A))=\frac{1}{N}\sum_{i=1}^{N}{\boldsymbol{z}^{(i)}}^{T}\log(A)\boldsymbol{z}^{(i)}=\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{v}^{(i)}}^{T}\log(A)\boldsymbol{v}^{(i)},$
@@
(3) $=\frac{n}{N}\sum_{i=1}^{N}\ {\boldsymbol{\mu}^{(i)}}^{T}\log(\Lambda)\boldsymbol{\mu}^{(i)}.$
Based on [18], the last term of equation (3) can be reformulated as a summation of Riemann Stieltjes integrals $I^{(i)}=\int_{a}^{b}\log(t)d\mu^{(i)}(t),i=1,\ldots,n$,
(4) $\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{\mu}^{(i)}}^{T}\log(\Lambda)\boldsymbol{\mu}^{(i)}$ $=\frac{n}{N}\sum_{i=1}^{N}\sum_{j=1}^{n}\log(\lambda_{j}^{(i)})\left[{\mu_{j}^{(i)}}\right]^{2}$
@@
(5) $\mu^{(i)}(t)=\left\{\begin{aligned} &\quad 0,&{\rm if}\ t<\lambda_{1}=a,\\ &\sum_{j=1}^{k-1}\left[{\mu_{j}^{(i)}}\right]^{2},&{\rm if}\ \lambda_{k-1}\leq t<\lambda_{k},k=2,...,n,\\ &\sum_{j=1}^{n}\left[{\mu_{j}^{(i)}}\right]^{2},&{\rm if}\ t\geq\lambda_{n}=b.\end{aligned}\right$.
@@
(6) $\frac{n}{N}\sum_{i=1}^{N}I^{(i)}\approx\frac{n}{N}\sum_{i=1}^{N}I_{m}^{(i)}=\frac{n}{N}\sum_{i=1}^{N}\sum_{k=1}^{m+1}\tau_{k}^{(i)}\log(\theta_{k}^{(i)}),$
where $m$ is the Lanczos step, and $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$, $\{\theta_{k}^{(i)}\}_{k=1}^{m+1}$ are the quadrature weights and nodes. According to [13], the weights $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$ are computed by the squares of the $1^{st}$ elements of the normalized eigenvectors $\{\boldsymbol{y}_{k}^{(i)}\}_{k=1}^{m+1}$ of $T_{m+1}^{(i)}$, and the nodes (also known as the Ritz values) $\{\theta_{k}^{(i)}\}_{k=1}^{m+1}$ are the corresponding eigenvalues $\{\lambda_{k}\}_{k=1}^{m+1}$ of $T_{m+1}^{(i)}$. Algorithm˜1 [12, Section 4.1] outlines how to compute $m+1$ Lanczos quadrature nodes and weights given a real symmetric matrix and a randomly generated Rademacher vector.
@@
10:Calculate the eigenpairs of $T_{m+1}$: $V,D]=\texttt{eig}(T_{m+1})$.
@@
Algorithm˜2 [33, Section 3] summarizes how the SLQ method approximates the log determinant for any SPD matrix.
@@
3:  Apply $m$-step Algorithm˜1 to AA with the initial vector $\boldsymbol{v}^{(i)}$ and obtain $\{\theta^{(i)}_{k}\}_{k=1}^{m+1}$ and $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$.
@@ 3 Error Analysis
According to [33, p. 9-10], the computational cost of stochastic Lanczos quadrature is $\mathcal{O}((\#nz(A)+nm^{2})N)$, where #nz(A)\#nz(A) denotes the number of non-zeros in symmetric matrix $A$ and $\mathcal{O}(nm^{2})$ represents the orthogonalization cost spent in the Lanczos iteration. Normally, the steps of Lanczos iteration $m$ and the number of Rademacher distributed vectors $N$ are expected to be much smaller than the matrix size $n$.
@@ 3.1 Error for symmetric quadrature nodes
[33, Theorem 4.2] Let $f$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci $\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $(m+1)$-point Lanczos quadrature approximation satisfies
(7) $|I-I_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-2}}\rho^{-2m-2}.$
@@
See [33, Section 4.1]. It should be pointed out that the result of Theorem˜3.1 assumes that the quadrature nodes are symmetric.
@@
[33, Corollary 4.5] Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $\lambda_{\min},\lambda_{\max}]$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
@@ 3.2 Existence of symmetric quadrature nodes
As previously mentioned, Theorem˜3.1 and Corollary˜3.3 are valid when the Gauss quadrature nodes are symmetric, but when do these nodes preserve this property? In order to answer this question, we propose the following Theorem˜3.6 with the aid of Definition˜3.5.
@@
(8) $A=\begin{bmatrix}A_{1,1}&\cdots&A_{1,q}\\ \vdots&&\vdots\\ A_{q,1}&\cdots&A_{q,q}\end{bmatrix}$,
@@
(9) $\begin{bmatrix}D_{1}&F\\ E&D_{2}\end{bmatrix}$,
@@
(10) $C=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}$,
where $O$ denotes matrix of zeros and $B\in\mathbb{C}^{m\times n}$. Such matrices have symmetric eigenvalue distribution about 0, proved by Theorem˜A.1 in Appendix˜A [21, 3]. Here we give a more concise theorem that proves the symmetry of eigenvalues when $m=n$.
@@
[…]
@@
Now we have checked the existence of matrices with symmetric eigenvalue distribution. Furthermore, when the conditions in Theorem˜3.6 hold, the Jacobi matrix $T_{m+1}$ generated by $m$-step Lanczos iteration has constant diagonal entries.
@@
(11) $\begin{bmatrix}\bar{\lambda}I_{m}&B\\ B^{H}&\bar{\lambda}I_{n}\end{bmatrix}$.
Thus, based on Corollary˜A.3, the Jacobi matrix $T_{m+1}$ generated by $m$-step Lanczos iteration have symmetric eigenvalue distribution, indicating that the $(m+1)$ quadrature nodes are symmetric about $\bar{\lambda}$. In addition, one may find several tests on Theorem˜3.6 in Appendix˜B.
@@ 3.3 Error for asymmetric quadrature nodes
Though Theorem˜3.6 gives sufficient conditions for symmetric quadrature nodes, the conditions are not always satisfied in practice. As shown in Section˜1, Fig.˜1 demonstrates that Lanczos quadrature nodes might be asymmetric. Therefore, Theorem˜3.12 derives an error estimation for the Lanczos quadrature with asymmetric quadrature nodes.
@@
[6, Corollary 3] Let $g$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci $\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $(m+1)$-point Lanczos quadrature approximation with asymmetric quadrature nodes satisfies
(12) $|I-I_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-1}}\rho^{-2m-2}.$
@@
The proofs of Theorem˜3.12 and Theorem˜3.1 differ by a single step, namely the calculation of an infinite geometric progression
(13) $|I-I_{m}|\leq\sum_{j=2m+2}^{\infty}\frac{4M_{\rho}}{\rho^{j}}.$
@@
(14) $\sum_{j=2m+2}^{\infty}\frac{4M_{\rho}}{\rho^{j}}=\frac{4M_{\rho}}{\rho^{2m+2}}\cdot\frac{1}{1-\rho^{-1}}=\frac{4M_{\rho}}{1-\rho^{-1}}\rho^{-2m-2}.$
@@
Theorem˜3.12 is in fact [6, Corollary 3] but with a redefinition of the parameter $m$. In [6, Algorithm 1], $m$ corresponds to $m$ quadrature nodes. Whereas, in Theorem˜3.12, $m$ corresponds to $(m+1)$ quadrature nodes, so does [33].
@@ 3.4 Error under affine transform
Theorem˜3.12 assumes that function $g$ is analytic in the reference domain $\left[-1,1\right]$ and then gives an error estimation in the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$. In order to conduct analysis back on the physical interval, a proper affine transform should be included. The SLQ method is based on a quadrature rule corresponding to a discontinuous measure (5). Unlike the continuous measure case, under such a discontinuous measure the quadrature error will remain unchanged when using an affine transformation from the reference domain $\left[-1,1\right]$ to the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$ [16]. Precisely, consider the integrals
$I^{(i)}=\int_{\lambda_{\min}}^{\lambda_{\max}}f(t)d\mu^{(i)}(t)$
@@
$J^{(i)}=\int_{-1}^{1}g(t)d\mu^{(i)}(h(t)),$
where $\mu^{(i)}$ are defined in (5), $h$ is an affine transform that shifts the interval [1,1][-1,1] to $\lambda_{\min},\lambda_{\max}]$, $g=f\circ h$, and $g$ satisfies the assumptions in Theorem˜3.12. These integrals are in fact the finite sums of the form
$J^{(i)}=\int_{-1}^{1}f(h(t))d\mu^{(i)}(h(t))$ $=\sum_{t_{j}:h(t_{j})=\lambda_{j}}f(h(t_{j}))\left[\mu^{(i)}(h(t_{j}))-\mu^{(i)}(h(t_{j})-)\right]$
@@
The measure in (15) does not satisfy assumption (16) since $d\mu^{(i)}(h(t))$ assigns positive measure $\left[\mu_{j}^{(i)}\right]^{2}$ on single point $h^{-1}(\lambda_{j})$, while the Lebesgue measure dtdt assigns zero on any single point. This may be the reason that (15) fails.